Two men and six boys can erect a fence in 12 day with the men working at the rate of one two- third of they boy how many men can erect the fence in 16 days.
Answer by ankor@dixie-net.com(22740)

Two men and six boys can erect a fence in 12 day with the men working at
the rate of one two-third of the boy
m = 5/3 boy
therefore:
b = 3/5 (one boy = .6 man)
;
Find how many man-days required to complete the job
12(2) + 12(6*.6) =
12(2) + 12(3.6) =
24 + 43.2 = 67.2 man-days
:
"how many men can erect the fence in 16 days."
16m = 67.2
m = 67.2/16
m = 4.2 men required to do the job in 16 days.